OpenStudy (anonymous):

write the equation of a line which passes through (5,-2) and is parallel to x=4?

OpenStudy (anonymous):

x= 5

OpenStudy (anonymous):

Parellel to x=4 means verticle lines, same x value through out so x=5

OpenStudy (anonymous):

can you show mw how to do it?

OpenStudy (anonymous):

x=4 is a straight line going up and down at where x=4. That means a parallel line would also be going straight up and down. For that up and down line to hit the point (5,-2) it would have to be at x=5

OpenStudy (anonymous):

x-4=0 is our line.. guess a equation of parellel for line x+k=0 it passes through (5,-2) hence 5+k=0 so k=-5 put it in our original equation x+k=0 hence answer is x-5=0

OpenStudy (anonymous):

uploading a picture in a sec if you want to visually see whats going on

OpenStudy (anonymous):